# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@
Abstract.
Large language models (LLMs) like OpenAI ChatGPT, Google Gemini, and GitHub Copilot are rapidly gaining traction in the software industry, but their full impact on software engineering remains insufficiently explored. Despite their growing adoption, there is a notable lack of formal, qualitative assessments of how LLMs are applied in real-world software development contexts. To fill this gap, we conducted semi-structured interviews with sixteen early-adopter professional developers to explore their use of LLMs throughout various stages of the software development life cycle. Our investigation examines four critical dimensions: people-how LLMs affect individual developers and teams; process-how LLMs alter software engineering workflows; product-LLM impact on software quality and innovation; and society-the broader socioeconomic and ethical implications of LLM adoption. Thematic analysis of our data reveals that while LLMs have not fundamentally revolutionized the development process, they have substantially enhanced routine coding tasks, including code generation, refactoring, and debugging. Developers who were LLM early-adopters report the most effective outcomes when providing LLMs with clear, well-defined problem statements, indicating that LLMs excel with decomposed problems and specific requirements. Furthermore, these early adopters identified that LLMs offer significant value for personal and professional development, aiding in the learning of new languages and concepts. Early adopters, highly skilled both in software engineering and in how LLMs work, identified early and persisting challenges for software engineering, such as inaccuracies in generated content and the need for careful manual review before integrating LLM outputs into production environments. Our study provides a nuanced understanding of how LLMs are shaping the current and future landscape of software development, highlighting both their practical benefits, limitations, and potential ongoing implications.
LLM, ChatGPT, Gemini, Copilot Chat, interview study, professional developers
journalyear: 2025ccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in HCI
1. Introduction
@@
Large language models (LLMs), trained on extensive datasets, are generative AI systems that produce human-like text based on their architecture and training data (OpenAI, 2023; Toner, 2023). Publicly-available LLMs such as OpenAI ChatGPT [^1] https://openai.com/chatgpt, Google Gemini [^2] https://gemini.google.com/, and GitHub Copilot Chat [^3] https://docs.github.com/en/copilot/github-copilot-chat represent transformative AI, reshaping workflows across multiple domains (Hoff et al., 2023; Mok and Zinkula, 2023; Verma and De Vynck, 2023). Within software development, specific LLMs are trained on vast repositories of code; for instance, the AI model behind GitHub Copilot was originally OpenAI’s Codex (n.a., 2021) and has since been refined to enhance contextual filtering (Zhao, 2024). These models enable developers to generate, refactor, and debug code through natural language prompts, streamlining traditionally labor-intensive processes. Such tools are increasingly central to enhancing developer productivity, fostering collaboration, and accelerating innovation in software engineering (Chen and Alias, 2024; Rane et al., 2023).
The adoption of LLMs in software development has sparked considerable discourse on their broader implications. For example, Stack Overflow’s annual developer surveys of 2023 and 2024 highlight how developers are rapidly integrating AI tools into their workflows, reflecting shifting perceptions about automation and its role in programming (n.a., 2023, 2024e) and researchers have studied. Discussions on LinkedIn (Augustin, 2023; (2024), DITS), the OpenAI community forum (Chen et al., 2024), and other platforms further emphasize the potential of LLMs to redefine software creation, engineering processes, and developer collaboration.
Several studies have examined LLM usage in software engineering through literature reviews (Hou et al., 2023), case studies (Belzner et al., 2023), empirical studies of forums (Chen et al., 2024), and comparison of general LLM tools and LLM-based agents (Jin et al., 2024). However, there is a lack of formal, qualitative assessments based on interviews with software developers on how LLMs are utilized in real-world software engineering settings. Those that exist primarily focus on more specific dimensions of their work, such as perceptions of security (Klemmer et al., 2024), trustworthiness (Rabani et al., 2023), and user study evaluations of new tools (Pinto et al., 2024).
@@
To address this gap, we conducted interviews with 16 early-adopter software industry professionals who took the initiative to become educated about LLMs at their inception and began actively incorporating LLM-based AI tools into their daily workflows between November 2022 and April 2023. The insights of these early adopters help frame new and ongoing issues that are important for software engineering processes, tools, and practitioners to maintain awareness of as LLMs evolve. Our investigation is organized around four critical dimensions: people, examining how LLMs influence individual developers and teams; process, analyzing changes in software engineering workflows; product, evaluating how LLMs contribute to software quality and innovation; and society, exploring the broader socioeconomic and ethical implications. By addressing these dimensions for both early and ongoing revisions of LLMs, this research provides a holistic perspective on the evolving intersection of LLMs and software development, offering insights into the opportunities and challenges these technologies bring to the field.
@@
    In this “people” dimension, we aimed to discern the advantages and disadvantages LLMs offer professional software developers in their work. While existing studies provide valuable insights into specific tasks where LLMs excel, they often fail to capture the broader, qualitative experiences of developers across varying experience levels. Our research fills this gap by exploring the nuanced ways LLMs influence different development tasks, offering a richer understanding of their practical utility and limitations. LLMs best serve developers when assisting with tasks they like the least, such as automating repetitive coding tasks and summarizing information. Additionally, LLMs helped developers learn more effectively by explaining code, personalizing learning, and generating new ideas. However, they are prone to hallucinations and inaccuracies, making it crucial for developers to manually review and adapt the generated content.
@@
    We explored the “process” dimension by investigating how LLMs have impacted the software development life cycle (SDLC), considering both positive and negative effects. Previous research has provided useful insights into isolated phases of the SDLC, but few studies examine the end-to-end impact of LLMs across all development stages. Our study fills this gap by systematically analyzing developers’ experiences using LLMs throughout the entire SDLC. We found that LLMs are particularly effective for ideation, testing, and debugging tasks, but less useful for generating requirements or reviewing code, especially in collaborative environments. Developers adapted their strategies by using a combination of broad and specific queries, experimenting with context addition and removal. While LLM-generated code often needed manual review before integration, it also offered learning opportunities and inspired innovative solutions.
@@
    In this “product” dimension, we focused on understanding the impact of LLMs on the code and software products generated by developers. Most existing studies rely on predefined metrics to evaluate code quality, often neglecting developers’ subjective perceptions of LLM-generated code. Our research addresses this gap by examining developers’ confidence in the code’s accuracy, readability, and complexity. We found that LLMs performed well with smaller, routine tasks like generating unit tests and documentation, but struggled with complex, novel code. Concerns arose around over-engineered code, security, and the sensitivity of the data used for training LLMs. Developers took responsibility for reviewing outputs, carefully balancing their trust in LLM-generated content with their own quality assurance processes.
@@
    To understand the broader societal implications, we explored developers’ perceptions of LLMs’ impact on the software industry and educational training. Much existing research focuses on theoretical implications, but few studies engage developers directly. Our research fills this gap by capturing their nuanced views on the opportunities and challenges of LLMs in professional workflows. While some of our participants noted that LLMs could replace certain roles, they generally believed developers’ roles were safe, as LLMs are tools to support rather than replace human decision-making. Concerns were raised about entry-level positions and the interview process, with some arguing for integrating LLMs into CS curricula, while others called for revising assignments to prevent easy LLM solutions. The lack of formal guidelines and training for LLM use in workplaces was noted, highlighting the need for structured support to maximize LLMs’ effectiveness and ethical use.
@@
  • We provide one of the first comprehensive qualitative analyses of developers’ experiences with LLM-based tools in real-world software development settings, offering new insights into their practical utility and limitations.
  • We uncover the strategies that developers use to integrate LLM tools into their workflows, highlighting both effective practices and common pitfalls, which can guide future adoption and optimization of these tools in professional environments.
  • We offer corroborating evidence for prior studies on the technical impact of LLMs on software development, while also introducing new insights into the broader socio-economic and educational implications of these tools, including their potential effects on job roles and educational curricula.
  • We analyze the impact of LLMs across all phases of the software development lifecycle (SDLC), providing a deeper understanding of how these tools influence technical tasks like coding, testing, and debugging, as well as human-centered activities like learning, collaboration, and decision-making.
The rest of the paper is structured as follows: Section 2 details the methodology used in the paper; Section 3 presents our results in detail for all four research questions; Section 4 explores the implications of our results; Section 5 discusses related work in the literature; Section 6 considers limitations; and Section 8 concludes the paper.
2. Methods
2.1. Participants and Demographics

### v2

@@
Abstract
Large language models (LLMs) are rapidly reshaping software development, but their impact across the full software development lifecycle is underexplored. Existing work tends to focus on isolated activities such as code generation or testing, leaving open questions about how LLMs affect developers, processes, products, and the broader software ecosystem.
We address this gap through semi-structured interviews with sixteen early-adopter software professionals who integrated LLM-based tools into their day-to-day work in early to mid-2023. We treat these interviews as early empirical evidence and, in the discussion, compare participants’ accounts with recent work on LLMs in software engineering, noting which early patterns persist or shift. Using thematic analysis, we organize our findings around four dimensions: people, process, product, and society. Developers reported substantial productivity gains from reducing mundane tasks, streamlining search, and accelerating debugging, but also described a productivity–quality paradox: they frequently discarded generated code and shifted effort from writing code to critically evaluating and integrating it. LLM use was highly phase-dependent, with strong uptake in implementation and debugging but limited influence on requirements gathering and collaborative work.
Participants developed new competencies to use LLMs effectively, including prompt engineering strategies, multi-layered verification, and security-conscious integration to protect proprietary data. They also anticipated changes in hiring expectations, team practices, and computing education, while emphasizing that human judgment and foundational software engineering skills remain essential. Our findings, later echoed in large-scale studies, offer actionable implications for developers, organizations, educators, and tool designers seeking to integrate LLMs responsibly into professional software practice.
keywords:
ChatGPT, Gemini, large language models, interview study, software engineering
1 Introduction
@@
Large language models (LLMs) are generative systems trained on extensive text and code corpora that can produce human-like responses to natural language prompts [1, 2]. Publicly-available LLMs such as OpenAI ChatGPT [^1] https://openai.com/chatgpt, Anthropic Claude [^2] https://claude.ai/, and Google Gemini [^3] https://gemini.google.com/ have rapidly reshaped workflows across multiple domains [3, 4, 5]. Within software development, LLMs trained on vast code repositories, such as OpenAI’s Codex [6, 7], enable developers to generate, refactor, and debug code through natural language prompts. These tools streamline formerly manual tasks and have become central to developer productivity, collaboration, and innovation in software engineering [8, 9].
The rapid adoption of LLMs in software development has prompted ongoing debate about how they should be integrated into professional practice. Stack Overflow’s annual developer surveys report growth from 44% to 76% of developers using or planning to use AI tools between 2023 and 2024 [10, 11], and enterprise deployments report substantial productivity gains [12]. However, concerns about code quality, security vulnerabilities, and impacts on developer learning have emerged alongside these benefits [13, 14].
Empirical work on LLMs in software engineering is still uneven. Existing studies predominantly address isolated activities, such as code generation [15], testing [16], or requirements engineering [17]. Other work has examined LLM usage in software engineering through literature reviews [hou2023large], case studies [18], empirical studies of forums [19], and comparison of general LLM tools and LLM-based agents [20]. Formal qualitative interviews with professional developers remain limited, and those that exist typically address narrow aspects such as security [21], trustworthiness [22], and user study evaluations of new tools [23], rather than tracing how LLMs impact people, processes, products, and organizational and societal issues across the software development lifecycle. Moreover, much of this empirical evidence was collected in the first wave of LLM adoption; assessing its continued relevance now requires explicitly relating early accounts to more recent studies.
@@
To address this gap, we conducted interviews with 16 early-adopter software industry professionals who took the initiative to become educated about LLMs at their inception and began actively incorporating LLM-based AI tools into their daily workflows between November 2022 and April 2023. These early adopters, interviewed in mid-2023, provide critical insights into emerging patterns that subsequent large-scale studies also report [24, 12, 11], which makes them particularly valuable for understanding how LLM integration has evolved in professional software development. A further objective of this study is to compare these early-adopter accounts with the rapidly expanding literature on LLMs in software engineering. We revisit this comparison in the discussion to illustrate how early patterns have persisted, evolved, or been challenged by subsequent work. Our analysis is organized around four dimensions that cover the software development lifecycle and its broader context: people, examining how LLMs influence individual developers and teams; process, analyzing changes in software engineering workflows; product, evaluating how LLMs contribute to software quality and innovation; and society, exploring the broader socioeconomic and ethical implications. This framework allows us to describe how LLMs have been reshaping software engineering practice, from individual task execution to organizational transformation.
@@
    We examine how LLMs change day-to-day development work, focusing on perceived benefits, risks, and shifts in skills. Our study’s participants reported clear gains in productivity by offloading mundane tasks, streamlining search, and utilizing boilerplate code. They also described LLMs as facilitating learning through personalizing education and explaining code, while also supporting personal growth by enhancing confidence and independence. However, developers reported significant challenges, including hallucinations and incomplete responses, limited knowledge for novel problems, struggles with logical reasoning, and concerns about impeding junior developer learning. These findings reveal a landscape where productivity gains must be balanced against quality assurance and educational concerns.
@@
    We mapped participants’ accounts onto the full software development lifecycle (SDLC), from requirements through maintenance. We found that LLMs have minimal impact on requirements gathering but are increasingly used for design ideation and problem decomposition. Implementation practices center on sophisticated prompt engineering strategies, where developers vary the specificity of prompts, use follow-up queries, and carefully manage security through generalized prompts. Code evaluation involves systematic verification through reading, output checking, and manual testing, with developers modifying code before integration and frequently discarding generated content. Developers turned to LLMs for generating unit tests and debugging, yet found them less reliable for complex tests, code review, and context window limitations. These findings demonstrate that LLM integration is highly phase-dependent, requiring developers to develop specialized strategies for different SDLC stages.
@@
    We investigated how developers perceive the impact of LLMs on the code and software products they create, focusing on accuracy, readability, complexity, and security. Participants generally viewed LLM-generated code as clean and readable for common, small tasks but raised concerns about over-engineered solutions, outdated information, and potential vulnerabilities. Although the majority of participants considered LLM code sufficiently secure for their non-production use cases, many expressed concerns about sending confidential data to LLMs and the need for careful review before using generated code in production. These findings highlight a quality paradox: LLMs can produce acceptable code for specific contexts but require substantial developer oversight to ensure data security and production readiness.
@@
    We explored developers’ perceptions of LLMs’ impact on the software industry and educational training. We found that developers are actively redefining LLMs’ organizational roles, comparing them to pair-programmers or junior developers, with some developers believing the overall job market will remain stable despite transformation. Widespread peer adoption coexists with concerns about entry-level positions and the need for organizational guidelines. In education, many participants advocated for integrating LLMs as learning tools, emphasizing the importance of teaching prompt engineering, while several stressed the continued importance of fundamental concepts and the need to adapt assignments and teaching methods. These views depict a profession in transition, grappling with how to use LLM capabilities while preserving core professional competencies and addressing equity concerns.
@@
  • We provide a qualitative analysis of developers’ experiences with LLM-based tools across the software development lifecycle, mapping reported impacts from requirements through maintenance and identifying phase-specific patterns of adoption and challenge.
  • We characterize the strategies developers use to integrate these tools into their workflows, including prompt engineering techniques, code verification practices, and context management approaches that can inform tool design and organizational guidelines.
  • We treat our 2023 interviews as early empirical evidence from early adopters and, in the discussion, compare these accounts with more recent empirical and conceptual work on LLMs in software engineering, clarifying where early patterns have remained stable, or evolved in the current literature.
  • We contribute empirical insights into the human factors and broader consequences of LLM adoption, including effects on developer confidence, learning, job roles, and educational practices, that complement technical evaluations and are critical for responsible use of LLMs in software engineering.
2 Related Works
The rapid integration of Large Language Models (LLMs) into software engineering practices necessitates a critical examination of their multifaceted impact. This paper provides such an examination across four key dimensions: the people (developers and teams), the processes (the development life cycle), the products (the software artifacts), and the broader societal implications. Due to thematic overlap, we have combined literature on the process and product subsections into a single section. The body of research in this domain is emerging at a significant pace. Although several systematic literature reviews have begun to map the field, such as recent work by Hou et al. on general LLM applications in SE [25] and Ferino et al. on novice developer perspectives [26], the sheer number of publications (both peer-reviewed and preprints) makes a comprehensive survey challenging. This aligns with observations by Fan et al. [27] that formal surveys struggle to capture all relevant, contemporary work. To reflect the current state of the field, we therefore include a small number of recent, formative non-peer-reviewed studies alongside peer-reviewed work.
2.1 People
2.1.1 Developer Perceptions and Well-being
Beyond broad systematic mappings, a growing body of qualitative work has begun to investigate developer-centric perspectives. Several interview studies of early adoption consistently report a tension between perceived productivity gains and concerns about tool reliability. For instance, Klemmer et al. found that while developers used AI assistants for security-critical tasks, they fundamentally lacked trust in the tool’s output and felt compelled to double-check its work [21]. This finding is mirrored in work by Rabani et al., who also noted developers’ emphasis on LLM inaccuracy and the subsequent need for manual debugging [22]. Similarly, Mendes et al. identified key challenges, including poor accuracy and the potential for distraction, which tempered the reported benefits of faster development [28]. While these studies capture developer perceptions, other work is beginning to quantify the impact of adoption on developer well-being. Feng et al., using a mixed-methods approach, modeled the relationship between Generative AI (GenAI) adoption and developer burnout, finding that GenAI adoption can heighten job demands and increase burnout, unless supported by positive attitudes and organizational resources [29].
2.1.2 Novice vs. Senior Developers
The “People” dimension of LLM adoption is further nuanced by developer experience, which strongly shapes the utility and risk of these tools. A body of work has concentrated on novice developers, for whom LLMs present a distinct dual-use scenario. A study by Kazemitabaar et al. indicates that while novices value LLMs for learning support, they concurrently struggle with over-reliance and face challenges in accurately evaluating AI-generated code [30]. This is reinforced by Siddiq et al., who observed that novices may treat LLMs as a “black box,” potentially impeding the development of fundamental skills [31]. In contrast, senior developers tend to employ LLMs as a “junior partner” for task augmentation, using them to generate boilerplate code or alternative implementations while relying on their own expertise to vet and refine the output [32].
2.1.3 Team Dynamics and Roles
Beyond the impact on individual developers, emerging research suggests that LLM integration is beginning to reconfigure team dynamics and collaborative processes. The acceleration of code generation, for instance, can introduce new pressures and bottlenecks in subsequent stages of the “Process,” particularly code review and quality assurance [33]. The “People” dimension is also highly role-dependent. Thapa et al., investigating the specific impact on software testers, found that LLMs are viewed with ambiguity compared to engineers: they are seen as valuable tools for augmenting test case generation but also as a potential threat that could devalue the human-centric and exploratory aspects of quality assurance [34].
2.2 Process and Product
2.2.1 Empirical Evaluation of LLMs in the SDLC
Moving from developer perceptions to empirical performance, another research vector has focused on evaluating LLMs against specific software engineering tasks. A primary focus has been on productivity and code generation. Peng et al., for example, conducted a controlled experiment observing a significant positive impact on developer productivity, with the AI-assisted group completing tasks more quickly [35]. Beyond general productivity, a growing number of studies are empirically assessing LLM capabilities in core “Process” and “Product” activities, particularly code refinement and quality assurance. For example, Guo et al. conducted an empirical study on the potential of ChatGPT for automated code refinement based on code reviews, finding that it outperformed a state-of-the-art baseline [36]. Similarly, Wadhwa et al. presented CORE, a tool utilizing an LLM-based proposer-ranker architecture to successfully resolve static code analysis warnings, demonstrating a practical application for improving code quality [37]. However, human oversight in these processes remains critical; both Alomar et al. and Siddiq et al. observed that although developers using AI-based code review tools were more efficient, they were also more likely to accept flawed code, highlighting a new set of risks for the “Process” dimension [38, 31].
2.2.2 Testing and Debugging
A significant sub-field of this empirical investigation concerns the application of LLMs to software testing and debugging—a core “Process” that directly governs “Product” quality. Zhang et al. present a comprehensive systematic literature review that concludes LLMs could automate the generation of a majority of unit tests and even identify potential bugs, albeit with varied performance [39]. Tanaka et al. further explored this by using LLMs to find and repair broken test cases, achieving a notable success rate [40]. More recently, research has focused on highly specific and challenging testing issues. Fatima et al., in their work on FlakyFix, successfully used LLMs not only to predict the category of fix required for a flaky test but also to perform the test code repair itself, demonstrating a guided repair process that leverages LLM capabilities [41]. These studies suggest a strong potential for LLMs to augment quality assurance workflows, although the reliability and scalability of these techniques are active areas of investigation.
2.2.3 Security Risks and Mitigation
The implications of LLMs for the software “Product” extend critically into the domain of security, presenting a significant dual-use challenge. On one hand, LLMs are being leveraged as a component of “Process” to improve security; for example, by powering advanced fuzzing frameworks [42] designed to detect vulnerabilities in complex contexts. On the other hand, a substantial body of work demonstrates that LLMs can inadvertently introduce critical vulnerabilities [43, 44]. This is compounded by the “self-deceive” problem, where an LLM’s programmatic goal to be helpful results in the generation of insecure code that developers, who may not be security experts, might overlook [45]. This risk has prompted a reactive line of research focused on mitigation, resulting in frameworks such as SafeGenBench, which aim to detect and mitigate security-related risks inherent in LLM-generated code [46].
2.2.4 Code Quality and Innovation
The integration of LLMs into the development “Process” has a direct, albeit complex, effect on the software “Product.” A primary area of concern is code quality. While LLMs can accelerate the quantity of code produced, a significant body of research, as previously discussed, has raised concerns about the quality and security of that output [46, 44, 43]. Beyond mere code generation, however, LLMs are also being positioned as tools for innovation and “Product” enhancement. For example, researchers are exploring their use in generating novel UI/UX designs and even in drafting initial software documentation, thereby expanding the scope of what constitutes the “Product” [47]. This suggests the impact on the “Product” is twofold: a high-volume, high-risk impact on code artifacts and a high-potential, nascent impact on innovation and non-code artifacts.
2.3 Society
2.3.1 Legal and Ethical Implications
Beyond the technical dimensions of ’People,’ ’Process,’ and ’Product,’ the adoption of LLMs introduces profound ’Societal’ implications, particularly regarding legal and ethical frameworks. The ambiguity surrounding the provenance of LLM training data has initiated a critical debate among legal scholars about copyright, data privacy, and intellectual property attribution [48]. These concerns are not merely theoretical; they have direct consequences for software development. Gabison et al. examined the practical legal challenges of liability for agentic agents and their outputs [44]. Furthermore, empirical studies such as LiCoEval have shown that LLM outputs may be subject to existing open-source licenses, creating complex and unresolved compliance issues for organizations that adopt these tools [49].
2.3.2 Impact on Computing Education
The integration of LLMs into pedagogical settings is being actively investigated to understand their effect on student learning and skill development. For example, Tanay et al. observed that upper-level computing students using LLMs for software engineering projects demonstrated improved efficiency in task completion and information retrieval [50]. Other work by Essel et al. similarly found that tools like ChatGPT could enhance undergraduate students’ critical, reflective, and creative thinking skills [51]. This emergent body of evidence has catalyzed discussions on curriculum adaptation. Ozkaya, for instance, has proposed that curricula must evolve to teach students how to collaborate effectively with AI, particularly for maintaining legacy systems [52]. Further supporting this, Jeuring et al. identified a strong correlation between computational thinking (CT) skills and a student’s ability to co-develop effectively with ChatGPT, reinforcing the continued necessity of teaching fundamental CT skills [53].
2.3.3 Literature Gap in Practitioner Perceived Effects on Education
While these studies provide valuable initial insights into how LLMs may augment student learning from an academic perspective, they illuminate a clear gap in the literature. There is, to date, limited research focused on the perspectives of professional developers regarding the role of LLMs in reshaping computing education. The views of active industry practitioners are a critical, missing component in the discussion of how to align educational practices with new, AI-driven professional settings. Our study helps to address this gap by engaging developers directly, capturing their experiences and views on how LLMs will influence the future teaching and learning of computing skills.
2.3.4 Perceptions and Risks
The “Societal” dimension of LLM adoption is characterized by a significant “double-edged sword” narrative, where perceived benefits are weighed against substantial risks. This duality is evident in public and developer discourse. On one hand, LLMs are seen as tools that can lower the barrier to entry for programming [54]. On the other hand, this same accessibility is a source of societal risk; several corporate threat assessment groups such as those at Google and Microsoft have begun reporting the use of LLMs for malicious purposes, such as generating malware or exploiting vulnerabilities at scale [55, 56]. This tension creates a complex “Societal” landscape where the technology is simultaneously viewed as a catalyst for innovation and a potential vector for significant harm.
2.3.5 Job Market and Displacement Anxiety
A dominant societal theme arising from widespread LLM adoption is the profound anxiety surrounding the future of the software engineering job market. The popular press and, to some extent, the academic literature have amplified concerns that LLMs could lead to widespread job displacement, de-skilling, or the devaluation of developer expertise [26, 25]. This narrative of automation-induced obsolescence is a frequent topic of investigation. However, research focusing on developer perceptions, including foundational work by Klemmer et al. and Mendes et al., suggests a more nuanced reality: developers often view these tools not as replacements, but as augmentations that shift, rather than eliminate, the need for human expertise [21, 28]. This perspective reframes the ”Societal” impact as an evolution of the developer role rather than its extinction.
Taken together, existing work offers rich but fragmented insights into how LLMs affect software engineering. Prior studies typically focus on specific roles (such as students, novices, testers, or individual developers), particular activities (such as code generation, testing, or security), or single dimensions of impact. There is still limited qualitative work that follows professional early adopters across the full software development lifecycle and situates their experiences within a broader sociotechnical framework spanning people, process, product, and society. Our study addresses this gap by providing an in-depth view of early LLM adoption in professional software development and by comparing these accounts with subsequent empirical and conceptual work to show how early patterns have persisted, or shifted.
3 Methods
3.1 Participants and Demographics